# Landscape of Thoughts: Visualizing the Reasoning Process of Large Language Models

### v1

@@ 6 Conclusion
This paper introduces the landscape of thoughts, a visualization tool for analyzing the reasoning paths produced by large language models with chain-of-thought. Built on top of feature vectors of intermediate states in reasoning paths, our tool reveals several insights into LLM reasoning, such as the relationship between convergence and accuracy, as well as issues like low consistency and high uncertainty. Our tool can also be adapted to predict the observed property, which is demonstrated by a lightweight verifier developed based on the feature vectors and our observations. We foresee that the landscape of thoughts will create several opportunities for developing, understanding, and monitoring LLM reasoning.
One limitation of the landscape of thoughts is its applicability only to multiple-choice tasks. Future work could focus on adapting this tool for open-ended reasoning tasks, such as mathematical problem-solving, code generation, and planning, where reasoning paths are less structured and more complex. Additionally, further research could aim to make the tool more accessible by generating intuitive visual and textual explanations, enabling non-experts to better understand and trust the reasoning processes of LLMs. Another promising direction is the development of automated methods to detect reasoning failures at scale, which could enhance the reliability of LLMs across diverse applications.
Impact Statement
Our work presents a tool for visualizing and understanding reasoning steps in large language models. We foresee our work will introduce more interpretability and transparency into the development and deployment of LLMs, advancing us toward more trustworthy machine learning. However, we must acknowledge that malicious activities can also be augmented by our tool. For example, attackers may use this tool to find prompts that bypass the alignment safeguards in LLMs. We believe such risks will be mitigated if this tool is widely adopted by safety researchers. Overall, the positive societal consequences of our work outweigh the negative ones, which stem primarily from misuse.
@@ Appendix A Further Discussions
Challenge 1: Bridging the token-thought gap. Current explanatory tools, including attention maps (Clark et al., 2019; Kobayashi et al., 2020), probing (Alain and Bengio, 2016; Tenney et al., 2019; Hewitt and Liang, 2019), and circuits (Elhage et al., 2021; Yao et al., 2024), primarily operate at the token-level explanation. While these approaches offer valuable insights into model inference, they struggle to capture the emergence of higher-level reasoning patterns from lower-level token interactions. Additionally, the discrete nature of natural language thoughts poses challenges for traditional statistical analysis tools designed for continuous spaces. Understanding how thought-level patterns contribute to complex reasoning capabilities requires new analytical frameworks that can bridge this conceptual gap.
Challenge 2: Analyzing without training data access. Existing investigations into LM reasoning have predominantly focused on correlating test questions with training data (Ippolito et al., 2022; Wang et al., 2024). This approach becomes particularly infeasible given the reality of modern LLMs: many models are closed-source, while some offer only model weights. Therefore, a desired analysis framework should operate across varying levels of model accessibility.
@@
Appendix B Supplementary Results and Analysis
B.1 Settings
@@
Visualizing the landscape of thoughts fundamentally relies on the decoding probability of LLMs. To this end, we adopted four open-source models with varying parameter sizes, namely Llama-3.2-1B, Llama-3.2-3B, Llama-3.1-8B, and Llama-3.1-70B. We repeatedly sample $10$ times from the target LLM using the same reasoning strategy as self-consistency (Wang et al., 2023b). For visualization purposes, we randomly sample $50$ questions from the testing split of each dataset and generate reasoning paths with the setup described above. For training the lightweight verifier, we randomly sample $20$ questions from the training split of each dataset to obtain the feature matrix ${\bm{S}}$. We extract these features using three model scales: Llama-3.2-3B, Llama-3.1-8B, and Llama-3.1-70B. Despite the relatively small training set, it proves sufficient for our lightweight verifier, which we subsequently evaluate on the data for visualization in Sec. 3.
B.2 Datasets
@@ B.3 Decoding Algorithms
Zero-shot CoT (Kojima et al., 2022). The core idea of this prompt strategy lies in adding simple instructions, e.g., ”Let’s think step by step.” to the prompt, enabling models to generate reasoning traces without assigned task-specific examples.

### v3

@@ 5 Conclusion
This paper introduces the landscape of thoughts, a visualization tool for analyzing the reasoning trajectories produced by large language models with chain-of-thought. Built on top of feature vectors of intermediate states in trajectories, our tool reveals several insights into LLM reasoning, such as the relationship between convergence and accuracy, and issues of low consistency and high uncertainty. Our tool can also be adapted to predict the answer of reasoning trajectories based on the observed property, which is demonstrated by a lightweight verifier developed based on the feature vectors and our observations for distinguishing the correctness of trajectories. We foresee that this tool will create several opportunities to develop, understand, and monitor the LLM reasoning.
Appendix
@@ A.1 Challenges in Analyzing LLM’s Reasoning Automatically
Challenge 1: Bridging the token-thought gap. Current explanatory tools, including attention maps (Clark et al., 2019; Kobayashi et al., 2020), probing (Alain & Bengio, 2016; Tenney et al., 2019; Hewitt & Liang, 2019), and circuits (Elhage et al., 2021; Yao et al., 2024), primarily operate at the token-level explanation. While these approaches offer valuable insights into model inference, they struggle to capture the emergence of higher-level reasoning patterns from lower-level token interactions. Additionally, the discrete nature of natural language thoughts poses challenges for traditional statistical analysis tools designed for continuous spaces. Understanding how thought-level patterns contribute to complex reasoning capabilities requires new analytical frameworks that can bridge this conceptual gap.
Challenge 2: Analyzing without training data access. Existing investigations into LM reasoning have predominantly focused on correlating test questions with training data (Ippolito et al., 2022; Wang et al., 2024a). This approach becomes particularly infeasible given the reality of modern LLMs: many models are closed-source, while some offer only model weights. Therefore, a desired analysis framework should operate across varying levels of model accessibility.
@@
A.2 A Comparison Between Landscape Visualization and Textual Analysis
Notably, for the language model, one could manually examine the responses to individual questions, as their responses are interpretable by humans. However, this approach has two major limitations:
Limitation 1: Lack of Scalability. Analyzing the individual question is time-consuming and labor-intensive. In general, text-based analysis requires human evaluators to carefully read long reasoning chains word by word. For example, if it takes 30 seconds to understand a single problem, reviewing 100 problems would require around 50 minutes of focused human effort. This burden grows quickly, especially as researchers often repeat this process many times while developing models and methods. In practice, researchers need quick, easily interpretable feedback, such as accuracy, when experimenting with changes to models and methods.
Limitation 2: Lack of Aggregation. It is difficult to aggregate insights across multiple problems to understand model behavior at the dataset level. Summarizing model behavior across multiple problems presents another challenge. Suppose one researcher has 100 reasoning chains; it is hard for him/her to reliably synthesize the model’s overall behavior. Different researchers may arrive at different, subjective summaries, which hinders consistency and interpretability.
By contrast, our visualization method provides a more objective and automatic way to analyze a model, making it much easier for researchers to analyze the model’s reasoning behavior. Similar to the t-SNE (van der Maaten & Hinton, 2008), the visualization enables a more comprehensive analysis of multiple reasoning problems instead of only one problem. The visualization uniquely combines human-readable paths with quantitative, scalable metrics for reasoning process analysis, enabling both model comparisons and mechanistic insights beyond manual text inspection.
Notably, the landscape provides unique insights into LLM reasoning that text analysis alone cannot capture. This power source bridges the gap between localized text understanding and global reasoning behavior. Our analysis in Sec. 3 reveals insights that are not revealed by previous text-based analysis. These insights include structural patterns across many reasoning paths, a strong correlation between early consistency and accuracy, and model-level differences where larger models explore more broadly than smaller ones.
A.3 The Intrinsic Relationship Between Visualization and Metrics
In the modeling of this work, we project each thought (state in a trajectory) from text space to numerical space, with the thought’s feature vector that each dimension indicates the distance to a particular answer (see Eqn. 1). We compute the feature vectors of all the thoughts from multiple trajectories and then obtain the feature matrix $\bm{F}$. Then, based on this feature matrix, we compute (1) the landscape visualization through dimension reduction and (2) the metrics of consistency and uncertainty. From this view, the metrics’ information can actually be seen from the landscape. In this work, we mainly focus on the landscapes and also use the metrics plots to help analyze.
In addition, landscape visualizations preserve the information of metrics, including the consistency, uncertainty, convergence, and many other metrics that are not covered in this work. The landscape provides a “global” view of the overall reasoning trajectories, while each metric provides a “local” view of a particular aspect. Note that humans naturally prefer visual matters like figures and videos, e.g., researchers prefer to use t-SNE in understanding the classification models. We recommend using landscape as a visualization tool to help understand the LLM reasoning, while the metric plots can further help inspect some particular aspects.
A.4 Discussion on Results and Observations
In the landscape visualizations, red regions map out the reasoning trajectories that end in incorrect answers, while blue regions map out those that end correctly. The contour lines and the depth of color together convey the density of reasoning states at each step: darker shades mean more trajectories passing through that region. As you observe a landscape evolve from its initial scatter of states toward later clustering, you’re seeing whether and how quickly the model’s reasoning paths lock onto an answer.
Observation 3.6 arises when we compare only the blue (correct) landscapes of different methods in Fig. 5. Early in the process, all methods scatter widely, exploring many possibilities; over time, though, some methods’ contours tighten more rapidly than others. Here, the landscape in Fig. 5 converges to its correct region much sooner—and with a denser cluster—than the landscapes in Fig. 5 to 5, and this faster, tighter convergence corresponds to its higher accuracy. Namely, methods with more scattered landscapes (converge more slowly) present lower accuracy than those that converge faster.
A related pattern appears when we compare models of different sizes in Fig. 2 (Observation 3.1). As we scale from the 1B model to the 70B model, the last 20% of the reasoning steps show increasingly dense blue clusters. Larger models, with greater capacity to store and retrieve information, steer their reasoning more directly and confidently toward the right answer, mirroring their higher accuracy. This further supports the positive correlation between convergence speed (of correct landscapes) and reasoning accuracy, which is revealed in Observation 3.6.
Observation 3.7 emerges from contrasting the red and blue landscapes of the same algorithm in Fig. 5. Here, failure trajectories (red) often settle into a wrong answer by roughly 20-40% of the reasoning process, while success trajectories (blue) only coalesce around the correct answer toward the very end—around 80-100% of the states. This indicates that early reasoning states are exploratory and can drift toward incorrect conclusions, whereas correct solutions only converge late in the trajectories. This convergence‐speed disparity between red and blue landscapes also holds across multiple datasets in Fig. 4.
Finally, Fig. 4 shows that each reasoning task leaves a distinct landscape “fingerprint,” supporting Observation 3.4. In AQuA, MMLU, and StrategyQA, the landscapes trace wide, structured sweeps of reasoning states—clear evidence of step-by-step deduction and exploration of intermediate hypotheses. By contrast, CommonSenseQA produces a tightly clustered trajectory from the outset, indicating direct retrieval of knowledge rather than an iterative trajectory. This divergence mirrors the tasks themselves: AQuA, MMLU, and StrategyQA require exploratory traversal through multiple reasoning steps, resulting in diverse yet organized state distributions, whereas CommonSenseQA depends on straightforward recall. These task-specific structures demonstrate how our landscape visualizations can uncover both shared patterns and fundamental differences across reasoning challenges.
In addition, each of these qualitative observations is further supported by statistical analyses in Appendix D.1, and we provide full visualizations, including annotated state trajectories (Figs. 23 to 26) and additional model comparisons (Figs. 27 to 28).
A.5 Potential Extension to Pruning Unpromising Trajectories
We showcase that our tool can be utilized to identify potentially incorrect reasoning trajectories at test time. In Section 3.3, we build up a lightweight verifier, which is based on the thoughts’ feature vectors and the consistency metric from the landscape of thoughts. This verifier indeed aims to predict the correctness of a reasoning trajectory, in order to boost the reasoning accuracy at test time. It is proven to be beneficial to the voting of multiple reasoning trajectories, as shown in Sec. 4.2.
Further, this verifier (together with the visualization tool) can be adopted to prune unpromising reasoning trajectories in tree-based searching. For instance, in methods like tree-of-thoughts and MCTS, a model explores multiple reasoning trajectories and usually uses the same model to identify the promising paths to search for the ultimate solution. Here, by leveraging features from the landscape of thoughts and the consistency metric, our verifier can identify flawed trajectories early during reasoning, acting as an efficient pruning mechanism to boost the search efficiency and reasoning performance.
Therefore, our tool can be integrated into the reasoning methods to monitor particular reasoning patterns (e.g., the correctness) and help understand as well as boost reasoning. There are multiple directions that deserve future exploration, including the one to identify and prune the potentially incorrect reasoning trajectories.
A.6 Potential Extension to Identify Post-hoc Trajectories
In the following, we discuss the feasibility of detecting post-hoc trajectory using our framework, particularly in defining the post-hoc trajectory. A post-hoc trajectory refers to the trajectory that the model exhibits high confidence in a single answer in the early states and maintains high consistency across states in the trajectory. Specifically,
  • the “early state” correspond to the “very early tokens of the response”;
  • the “high confidence in a single answer” corresponds to the “model has chosen its answer“;
  • the “high consistency across states in the trajectory” corresponds to the “trajectory is produced as a consequence of that decision”.
Namely, the post-hoc trajectory can be potentially identified by inspecting the confidence and consistency of particular positions of states in our framework. Then, we elaborate on the more detailed definitions for the three components above.
  • For defining the “early states”, it should have an absolute threshold of states index, e.g., early 10 states, or a relative threshold, e.g., early 10% of states. This threshold should be chosen deliberately, and the states with an index smaller than this threshold are categorized as “early states”.
  • Similarly, a clear threshold is necessary for defining the “high confidence” or “high consistency”, e.g., over 80% confidence and 60% consistency. With the metrics defined in Section 2.3, here, we should examine (1) the confidence of the early states in the trajectory and (2) the consistency across all states of the trajectory. Here, only the trajectory that exceeds the confidence threshold as well as the consistency threshold can be classified as a post-hoc trajectory.
In conclusion, our framework shows promise for identifying post-hoc trajectories. Meanwhile, we should note that it still needs (1) to choose particular thresholds for the precise definition of post-hoc trajectory and (2) to collect a set of reliable data to verify the effectiveness in identifying post-hoc trajectory. These are quite challenging to conduct. Although it goes beyond the scope of work, we believe investigating post-hoc trajectory in reasoning is valuable and merits exploration in future work.
A.7 Limitations and Future Directions
Scope. While the Landscape of Thoughts offers a practical lens on model reasoning, its current instantiation is limited to multiple-choice settings. Extending LoT to open-ended reasoning—including mathematical problem solving, code generation, and planning—requires handling less structured and more entangled reasoning paths. Two complementary threads of future work are: (i) improving accessibility by producing intuitive visual and textual explanations that help non-experts inspect and trust model behavior, and (ii) developing automated, scalable detectors of reasoning failures to improve reliability across applications.
Key challenge: synthesizing options. The central obstacle is the quality of the synthesized answer options. Human-authored distractors are carefully calibrated to be plausible, exposing distinctions between (1) correct reasoning and (2) reasonable-but-wrong reasoning (e.g., overlooking information or making arithmetic slips). In contrast, LLM-generated distractors can be implausible and thus trivially eliminated when juxtaposed with the correct option, yielding visualizations that over-emphasize the correct trace and limit diagnostic value. Moreover, LLMs may reuse similar reasoning patterns, producing near-duplicate error modes across incorrect options and reducing the comprehensiveness of the analysis.
Mitigations. To address these issues, we can elicit higher-quality distractors with state-of-the-art LLMs (e.g., OpenAI o3, Gemini 2.5 Pro) and tune sampling hyperparameters (temperature, top-$p$) to promote diversity and explore alternative solution trajectories.
Binary reformulation. A practical alternative is to recast multiple-choice prompts as binary (yes/no) queries. For example, the question “What is the capital of France?” can be reformulated as “Is Paris the capital of France?” with options Yes or No. Under this framing, both options remain prima facie plausible: the incorrect choice admits coherent yet flawed rationales, and the variety of “No” trajectories preserves diversity without resorting to obviously implausible distractors.
Beyond multiple choice. Although open-ended tasks are beyond the present scope, LoT is, in principle, extendable. The key requirement is to construct a candidate set of answers by querying the model (a non-trivial step that is given for free in multiple-choice tasks). Treat the ground-truth answer as one option and generate additional plausible alternatives using LLMs; LoT can then analyze the induced reasoning behaviors in these open-ended scenarios.
Case: code generation. Code generation introduces additional challenges: there is typically no single ground-truth program, and evaluation proceeds via test suites. Candidate programs are diverse and do not naturally discretize into options. We propose the following procedure: (i) sample multiple candidate solutions from the model under evaluation; (ii) score each by the number of tests passed; (iii) apply a threshold to separate more-correct from less-correct solutions; (iv) embed and cluster solutions within each partition; and (v) use cluster centroids as anchors for “correct” and “incorrect” choices. Cluster quality can be assessed with the Silhouette Score and the Davies-Bouldin Index. These anchors enable a LoT-style visualization over the solution space and provide insight into reasoning behaviors.
In summary, our visualization framework is adaptable beyond multiple-choice scenarios. To our knowledge, LoT is the first landscape visualization tool aimed at analyzing LLM reasoning; it is imperfect and remains open to improvement and extension. We believe it constitutes a small but meaningful step toward understanding and improving the reasoning processes of LLMs.
A.8 A Comparison Between Lightweight Verifier and Reward-guided Algorithms
It is worth noting that our goal is not to build a sophisticated verifier, but rather to demonstrate how the feature vectors from the landscape visualization can be effectively used.
In general, reward-guided algorithms are more computationally efficient than the path landscape. Specifically, for a reasoning path with $n$ thoughts and $c$ answer choices, constructing the landscape requires $n\times c$ forward passes through the reasoning model. In contrast, a reward-guided approach typically makes a single call to a reward model that evaluates the entire reasoning chain at once.
Meanwhile, it’s important to consider the overhead involved in training the reward models in reward-guided algorithms. Notably, for Process-Reward Models (PRMs) (Luo et al., 2024; Xu et al., 2025), collecting high-quality training data often requires detailed, fine-grained annotations of reasoning steps, which can be costly and time-consuming. Moreover, training a reward model (often itself an LLM) incurs significant computational expense. In contrast, our lightweight verifier is much more efficient to train, as it requires no human annotations and uses easily obtainable data.
Appendix B Related Work
Reasoning with large language models. Chain-of-Thought (CoT) prompting (Wei et al., 2022; Kojima et al., 2022) has empowered LLMs to tackle multi-step reasoning problems by generating intermediate steps before producing a final answer. Building upon CoT, numerous methods have been proposed to address various challenges, including compositional generalization (Zhou et al., 2023; Khot et al., 2023), planning (Yao et al., 2023a; Hao et al., 2023), and rule learning (Zhu et al., 2023) within the CoT reasoning. Beyond solving reasoning tasks, CoT has also emerged as a foundational framework for other techniques, such as fine-tuning LLMs (Zelikman et al., 2022), enabling LLM-based agents (Yao et al., 2023b), and facilitating test-time scaling (Snell et al., 2024). Nevertheless, most of these approaches are developed in a trial-and-error manner, largely due to the absence of proper tools for analyzing the CoT.
Understanding chain-of-thought reasoning. There are a few studies that explore what makes CoT prompting effective by perturbing its exemplars. To be specific, Madaan & Yazdanbakhsh (2022) found that the text and patterns of exemplars help CoT generate sentences resembling correct answers. Besides, Wang et al. (2023a) highlighted the importance of maintaining the correct order of reasoning steps, while Ye et al. (2022) demonstrated that using complementary exemplars can enhance reasoning performance. Furthermore, CoT can benefit from longer reasoning chains, even without new information to the prompt (Jin et al., 2024). Another line of research investigates CoT’s general behavior (Tang et al., 2023; Saparov & He, 2023; Saparov et al., 2023; Shi et al., 2023). For example, CoT heavily depends on the semantic structure of the problem to perform reasoning (Tang et al., 2023), struggles with planning and unification in deductive reasoning (Saparov & He, 2023), has difficulty generalizing to longer reasoning paths (Saparov et al., 2023), and can be easily misled by irrelevant information in the context (Shi et al., 2023). However, these observations are derived from specific reasoning tasks and prompt settings, limiting their applicability to other scenarios. In contrast, we introduce a general-purpose tool that allows users to analyze reasoning in their contexts.
Tools for analyzing chain-of-thought. To the best of our knowledge, the only existing tool for analyzing CoT is gradient-based feature attribution (Wu et al., 2023), which computes a saliency score for each input token based on the model’s output. However, these token-level saliency scores do not directly capture the thought-level, multi-step reasoning process of LLMs. Consequently, the main finding in (Wu et al., 2023) is that CoT stabilizes saliency scores on semantically relevant tokens compared to direct prompting. Metrics designed to quantify CoT performance (Chen et al., 2024; Ton et al., 2024) can also be used to analyze the reasoning behaviors of LLMs. For instance, Ton et al. (2024) employs information gain to identify failure modes in reasoning paths, aligning with Observation 3.7 in this paper. However, our 2-D visualization offers significantly deeper insights than a single information gain metric. Additionally, the verifier derived from our tool is conceptually related to outcome-supervised reward models (Cobbe et al., 2021).
Measuring uncertainty and consistency in LLM reasoning. Several works in this research line compute metrics (such as confidence and perplexity) by leveraging the features from LLMs to measure and detect hallucination in reasoning (Li et al., 2023; Chuang et al., 2024; Yang et al., 2025). Specifically, low confidence and high perplexity often indicate unreliable reasoning, enabling the development of lightweight detectors to guide reasoning and mitigate hallucinations. However, these metrics have limitations (Xiong et al., 2024; Zhang et al., 2023): they can exhibit over-confidence or low perplexity in incorrect responses, their reliability relies heavily on the models’ capability, and they cannot provide more comprehensive insights into the multiple reasoning trajectories. By contrast, our landscape of thoughts offers a holistic approach, integrating several existing metrics. This framework enables global qualitative analysis, including measures of perplexity, consistency, and uncertainty. In addition, the landscape of thoughts enables the development of advanced tools to enhance reasoning by using the features and metrics, as mentioned in Sec. 3.3.
Appendix C Experiment Settings
C.1 Setup
@@
Visualizing the landscape of thoughts fundamentally relies on the decoding probability of LLMs. To this end, we adopted four open-source models with varying parameter sizes, namely Llama-3.2-1B, Llama-3.2-3B, Llama-3.1-8B, and Llama-3.1-70B. We repeatedly sample $10$ times from the target LLM using the same reasoning strategy as self-consistency (Wang et al., 2023b).
For visualization purposes, we randomly sample $50$ questions from the testing split of each dataset and generate reasoning paths with the setup described above. For simplicity, we compute distances only between each state and all candidate answers. To visualize multiple problems in a shared space, we always place the distance to the correct answer as the first element of each feature vector. This alignment allows joint analysis across problems, as introduced in the paragraph below Equation 4. We then aggregate feature vectors from all problems into a feature matrix (Equation 2), which is passed to t-SNE to compute the pairwise distance between any two states and then outputs the 2D coordinate of each state.
For training the lightweight verifier, we randomly sample $20$ questions from the training split of each dataset to obtain the feature matrix $\bm{S}$}. We extract these features using three model scales: Llama-3.2-3B, Llama-3.1-8B, and Llama-3.1-70B. Despite the relatively small training set, it proves sufficient for our lightweight verifier, which we subsequently evaluate on the data for visualization in Sec. 3.
C.2 Datasets
@@ C.3 Decoding Algorithms
Zero-shot CoT (Kojima et al., 2022). The core idea of this prompt strategy lies in adding simple instructions, e.g., "Let’s think step by step." to the prompt, enabling models to generate reasoning traces without assigned task-specific examples.